OpenStudy (anonymous):

I need help answering questions like this one a[n]=2 a[n-1]-a[n-2] ; a[1]=4 ; a[2]=3

OpenStudy (anonymous):

a[n] is equivalence classes ?

OpenStudy (anonymous):

no its a Subscript

OpenStudy (anonymous):

this is supposed to be a geometric sequence

OpenStudy (anonymous):

Whats the question?

OpenStudy (anonymous):

for 3rd term a[3] = 2*a[3-1] - a[3-2] = 2*a[2] - a[1] = 3 - 4 = -1

OpenStudy (math&ing001):

@Ignaris to find the other terms of the sequence just replace n with the number of the sequence. For example for 3: a[3] = 2*a[3-1] - a[3-2] = 2*a[2] - a[1] = 2*3 - 4 = 6 - 4 = 2 @sangya21 is right, he just made a typo.

OpenStudy (anonymous):

Thanks Math&ing001 :)

OpenStudy (math&ing001):

No probs =)

OpenStudy (anonymous):

I kept trying to put in the calculator and got the wrong answer X)

OpenStudy (math&ing001):

Well you don't really need a calculator for this one =p